# The General Position Problem in Graph Theory: A Survey

### v3

@@ 1 Introduction
[…]
@@
No-three-in-line problems are also familiar from games that span multiple cultures from the start of recorded history to the modern day. The ancient Egyptians and Romans played games in which the winner had to get three in a line [128], leading up to our modern day games Tic-Tac-Toe (known to the third author in the UK as ‘Noughts and Crosses’) and ‘Nine Men’s Morris’ (also known as ‘Mill’). For Tic-Tac-Toe and Nine Men’s Morris, optimal play leads to a draw [50], i.e. a configuration with no three counters in line to which no further counters can be added. The game ‘Connect Four’ requires a player to get four counters in a line to win, whereas in the Japanese game Gomoku the winner must achieve five in a row (the former game is a first player win [5] and the latter is a first player win on a $\times 15$ 15 board [6]). We could mention many other examples. Some mathematical discussion of these games can be found in [59].
@@
The mathematical investigation of such problems in combinatorics can be dated to a puzzle of the famous recreational mathematician Henry Dudeney (1857-1930). As a keen chess player, many of his puzzles are set on a chessboard, and one such problem from 1900 (see [40]) can be regarded as the origin of the general position problem. The puzzle asks for the largest number of pawns that can be placed on an $\times 8$ 8 chessboard without three pawns lying on any straight line in the plane (not just rows, columns or diagonals of the board). The answer is 16. The more general No-Three-In-Line Problem for an $n\times n$ chessboard has been called by Brass, Moser and Pach “one of the oldest and most extensively studied geometric questions concerning lattice points” [15].
[…]
@@
For a lower bound on the $n\times n$ board, Erdős gave a construction containing $n-o(n)$ pawns in general position, published by Roth in [102] (interestingly Erdős investigated this as a way to attack the Heilbronn triangle problem). This estimate was later improved to $\frac{3n}{2}-o(n)$ pawns by Hall et al. [55]. A probabilistic argument by Guy and Kelly [53, 54] (with a subsequent slight correction by Guy) suggests that the true answer is $\frac{\pi n}{\sqrt{3}}+O(n)$, although see [52] for the summary of a talk by Kaplan that gives a counterargument.
@@
The No-Three-In-Line Problem has been investigated in other geometrical settings. The article [86] examines the density of No-Three-In-Line sets on the infinite grid; they give a construction that contains $\Theta\left(\frac{n}{\log^{1+\epsilon}n}\right)$ vertices from $[1,n]^{2}$ for every $n$ and conjecture that this can be improved to a linear lower bound. The paper [95] treats the No-Three-In-Line Problem on a three-dimensional grid and a higher dimensional version can be found in [108], whereas the papers [31, 78, 85] treat the problem on a torus. The number of ways to arrange $k$ points in an $n\times n\times n$ triangular grid is given in OEIS sequence A194136. Other geometrical investigations into general position sets include [9, 27, 42]. The general position subset selection problem in discrete geometry also has an extensive literature, including [48, 93], of which we will not attempt to give an exhaustive overview.
Figure 1: A configuration of 104 pawns in general position on a $\times 52$ 52 chessboard [46]
@@ 1.1 Terminology
For a positive integer $k$ we will use the notation $[k]=\{1,\ldots,k\}$. We will denote the order and the size of a graph $G$ by $n(G)$ and $m(G)$ respectively. If $u\in V(G)$, then $N_{G}(u)$ denotes the open neighbourhood of $u$ and $N_{G}[u]$ its closed neighbourhood. Vertices $u$ and $v$ of a graph $G$ are true twins if $N_{G}[u]=N_{G}[v]$. A vertex is a support vertex if it is adjacent to a leaf. A vertex $u$ of $G$ is simplicial if its neighbourhood induces a clique. The set of all simplicial vertices of a graph will be denoted by $S(G)$ and their number by $\operatorname{s}(G)$. The number of leaves of a tree $T$ will be written as $\ell(T)$ (so $\ell(T)=\operatorname{s}(T)$). The join of $G$ and $H$ is denoted by $G\vee H$. For $X\subseteq V(G)$, $G[X]$ is the subgraph of $G$ induced by $X$. By $\alpha(G)$ and $\omega(G)$ we represent the independence number and the clique number of $G$ respectively.
@@
[…]
@@ 2 General Position Number
General position sets have a structure that can be characterised. To do this, we need the following definitions. Let $G$ be a connected graph, $X\subseteq V(G)$, and ${\cal P}=\{X_{1},\ldots,X_{p}\}$ a partition of $X$. Then ${\cal P}$ is distance-constant if for any $i,j\in[p]$, $i\neq j$, the distance $d_{G}(u,v)$, $u\in X_{i}$, $v\in X_{j}$, is independent of the selection of $u$ and $v$. If ${\cal P}$ is a distance-constant partition, and $i,j\in[p]$, $i\neq j$, then the distance $d_{G}(X_{i},X_{j})$ can be defined as the distance between a vertex of $X_{i}$ and a vertex of $X_{j}$. A distance-constant partition ${\cal P}$ is in-transitive if $d_{G}(X_{i},X_{k})\neq d_{G}(X_{i},X_{j})+d_{G}(X_{j},X_{k})$ holds for arbitrary pairwise distinct $i,j,k\in[p]$.
@@ 2.1 Bounds
Graphs with $\operatorname{gp}(G)\in\{2,n(G),n(G)-1\}$ are characterised in the paper [22]. In particular, $\operatorname{gp}(G)=n(G)$ if and only if $G=K_{n}$, and $\operatorname{gp}(G)=2$ if and only if $G\in\{C_{4},P_{n}\}$ [22, Theorem 2.10]. By [22, Theorem 2.4], for any connected graph $G$ we have $\operatorname{gp}(G)\leq n(G)-\mathrm{diam}(G)+1$. Consequently, [22, Theorem 3.1] proves that $\operatorname{gp}(G)=n(G)-1$ if and only if $G=K_{1}+\bigcup_{j}m_{j}K_{j}$ with $\sum m_{j}\geq 2$ or $G=K_{n}-\{e_{1},\ldots,e_{k}\}$ with $k\in[n-3]$, where the $e_{i}$’s are edges in $K_{n}$ that are all incident to a common vertex.
@@
[110, Theorem 3.1] If $G$ is a connected graph of order at least four, then ${\rm gp}(G)=n(G)-2$ if and only if $G\in\cup_{i=1}^{8}{\cal F}_{i}.$
@@
[80, Theorem 3.1] (Isometric Cover Lemma) If $\left\{H_{1},\ldots,H_{k}\right\}$ is an isometric cover of $G$, then
@@
The isometric-path number of a graph $G$, which we will write as $\rho(G)$, is the smallest number of isometric paths (shortest paths) needed to cover all the vertices of $G$ (note that $\operatorname{ip}(G)$ is often used for this invariant, but we wish to avoid a notation clash with the independent position number). Similarly, the isometric cycle number of $G$, written as $\rho_{C}(G)$, is the smallest number of isometric cycles required to cover all the vertices of $G$. If $G$ cannot admit a cover by isometric cycles, we set $\rho_{C}(G)=\infty$. Then, the Isometric Cover Lemma yields that for any graph $G$, $\operatorname{gp}(G)\leq 2\rho(G)$, and $\operatorname{gp}(G)\leq 3\rho_{C}(G)$ (see [80, Corollary 3.2]).
@@
[80, Theorem 3.3] If $R$ is a general position set of a graph $G$ and $v\in R$, then
@@
[51, Theorem 4.1] If $G$ is a connected graph, then $\operatorname{gp}(G)\leq\alpha^{\omega}(G)$. Moreover, if $\operatorname{diam}(G)\in[2]$, then $\operatorname{gp}(G)=\alpha^{\omega}(G)$.
@@ 2.1.1 Strong resolving graphs
[71, Theorem 3.1] If $G$ is a connected graph, then $\operatorname{gp}(G)\geq\omega(G\operatorname{{}_{SR}})$. Moreover, equality holds if and only if $G$ contains a $\operatorname{gp}$-set that induces a complete subgraph of $G\operatorname{{}_{SR}}$.
@@
[71, Proposition 2.4] If $G$ has no true twins and $\operatorname{diam}(G)=2$, then $\operatorname{gp}(G)=\omega(G\operatorname{{}_{SR}})$ if and only if $\operatorname{gp}(G)=\alpha(G)$.
On the other hand, as demonstrated in [71, Proposition 3.5], for any integers $r\geq t\geq 2$, there exists a graph $G$ such that $\operatorname{gp}(G)=r$ and $\omega(G\operatorname{{}_{SR}})=t$.
@@ 2.2 Complexity
[80, Theorem 5.1] General position set is NP-Complete.
@@ 2.3 Exact results
As the general position problem is NP-complete, it makes sense to consider the restriction of the problem to some simple graph classes. We recall from the seminal papers [22, 80] the following results. If $n\geq 2$, then $\operatorname{gp}(K_{n})=n$. If $n\geq 5$, then $\operatorname{gp}(C_{n})=3$, while $\operatorname{gp}(C_{4})=2$. If $n\geq 2$, then $\operatorname{gp}(P_{n})=2$. Further, if $G$ is a block graph, then $\operatorname{gp}(G)=\operatorname{s}(G)$ [80, Theorem 3.6]. Hence in particular, if $T$ is a tree, then $\operatorname{gp}(T)=\ell(T)$ [22, Theorem 2.5].
@@ 2.3.1 Kneser graphs
[…]
@@
[51, Theorem 2.3] Let $n$ and $k$ be positive integers with $n\geq 3k-1$. If, for all $t$ such that $2\leq t\leq k$, the inequality $k^{t}\binom{n-t}{k-t}+t\leq\binom{n-1}{k-1}$ holds, then $\operatorname{gp}(K(n,k))=\binom{n-1}{k-1}$.
For fixed $k$ and $t=2$, the above inequality is satisfied when $n\geq k^{3}-k^{2}+2k-1$, while Patkós [91, Theorem 1.2] improved this result by proving that the same conclusion holds for $n\geq 2.5k-0.5$. The result is stated as follows.
@@
[91, Theorem 1.2] If $n,k\geq 4$ are integers with $n\geq 2k+1$, then $\operatorname{gp}(K(n,k))\leq\binom{n-1}{k-1}$ holds. Moreover, if $n\geq 2.5k-0.5$, then we have $\operatorname{gp}(K(n,k))=\binom{n-1}{k-1}$, while if $2k+1\leq n<2.5k-0.5$, then $\operatorname{gp}(K(n,k))<\binom{n-1}{k-1}$ holds.

### v4

# The General Position Problem: A Survey

@@ 1 Introduction
Many significant mathematical problems involve several objects lying on a line, or else how to avoid such configurations. For example, a collection of vectors is in general position if no three lie on a common line, so that such problems relate to the fundamental notions of dimension and independence. The Hales-Jewett Theorem and other results from Ramsey theory on the integers and combinatorial set theory can be interpreted as no-$k$-in-line problems, and the study of points with no $k$ on a line has a long history in discrete geometry.
@@
No-three-in-line problems are also familiar from games that span multiple cultures from the start of recorded history to the modern day. The ancient Egyptians and Romans played games in which the winner had to get three in a line [146], leading up to our modern day games Tic-Tac-Toe (known to the third author in the UK as ‘Noughts and Crosses’) and ‘Nine Men’s Morris’ (also known as ‘Mill’). For Tic-Tac-Toe and Nine Men’s Morris, optimal play leads to a draw [55], i.e. a configuration with no three counters in line to which no further counters can be added. The game ‘Connect Four’ requires a player to get four counters in a line to win, whereas in the Japanese game Gomoku the winner must achieve five in a row (the former game is a first player win [5] and the latter is a first player win on a $15\times 15$ board [6]). We could mention many other examples. Some mathematical discussion of these games can be found in [67].
@@
The mathematical investigation of such problems in combinatorics can be dated to a puzzle of the famous recreational mathematician Henry Dudeney (1857-1930). As a keen chess player, many of his puzzles are set on a chessboard, and one such problem from 1900 (see [45]) can be regarded as the origin of the general position problem. The puzzle asks for the largest number of pawns that can be placed on an $8\times 8$ chessboard without three pawns lying on any straight line in the plane (not just rows, columns or diagonals of the board). The answer is 16. The more general No-Three-In-Line Problem for an $n\times n$ chessboard has been called by Brass, Moser and Pach “one of the oldest and most extensively studied geometric questions concerning lattice points” [17].
For the $n\times n$ board, since any row or column can contain at most two pawns, $2n$ pawns is a trivial upper bound. For many values of $n$ this upper bound can be achieved [50, 51, 60, 110, 111], including all $n$ in the range $2\leq n\leq 60$, with $n=61$ currently the smallest open case [110]. A solution for $n=60$ with 120 pawns for $n=60$ due to Prellberg is shown in Figure 1. 62,563 different configurations meeting the $2n$ upper bound can be found at [37]. There are typically many maximum configurations (the number of solutions with $2n$ pawns is given in OEIS entry A000769), and the aforementioned computational results are often concerned with finding maximum configurations with certain symmetry properties. Applications of the problem include optimal graph drawings with edges as straight line segments [16, 143].
@@
For a lower bound on the $n\times n$ board, Erdős gave a construction containing no(n)n-o(n) pawns in general position, published by Roth in [116] (interestingly Erdős investigated this as a way to attack the Heilbronn triangle problem). This estimate was later improved to $\frac{3n}{2}-o(n)$ pawns by Hall et al. [61]. A probabilistic argument by Guy and Kelly [59, 60] (with a subsequent slight correction by Guy) suggests that the true answer is $\frac{\pi n}{\sqrt{3}}+O(n)$, although see [57] for the summary of a talk by Kaplan that gives a counterargument. It is noted in [110] that the heuristic probabilistic argument only applies for $n\geq 493$, and so the fact that the $2n$ upper bound can be met for small $n$ is not strong evidence against the conjecture of Guy and Kelly. An early approach to the no-three-in-line problem using neural networks can be found in [138], whilst a recent approach using machine learning and AI is presented and contrasted with the results of integer linear programming in [112].
The more general problem of how to place pawns on the $n\times n$ grid without $k$-in-line was treated in recently in [85] and [86]. The paper [85] gives a randomised algebraic construction that places $\left(1-\frac{2}{k-1}\right)kn$ or $\left(1-\frac{3}{k-1}\right)kn$ pawns on the grid with no $k$-in-line for odd and even $k$ respectively, with better bounds for $k\leq 21$. The fact that the $2n$ upper bound for $k=2$ can be met for small $n$ is generalised in [86] and [58] to show that $kn$ pawns can be positioned without creating $(k+1)$-in-a-line for $k\geq\frac{5\sqrt{35}}{2}\sqrt{n\log n}$ and $k\geq 10^{37}$ respectively.
@@
The No-Three-In-Line Problem has been investigated in other geometrical settings. The article [97] examines the density of No-Three-In-Line sets on the infinite grid; they give a construction that contains $\Theta\left(\frac{n}{\log^{1+\epsilon}n}\right)$ vertices from $1,n]^{2}$ for every $n$ and conjecture that this can be improved to a linear lower bound. The paper [106] treats the No-Three-In-Line Problem on a three-dimensional grid and a higher dimensional version can be found in [123], whereas the papers [35, 89, 96] treat the problem on a torus. The two articles [44, 122] investigate the problem of finding large subsets of the $d$-dimensional grid with no $d+2$ vertices lying on a common hyperplane or $d$-dimensional sphere, with an extension to quadratic surfaces in [124]. A generalised version in which configurations other than straight lines are forbidden is discussed in [70]. The number of ways to arrange $k$ points in an $n\times n\times n$ triangular grid is given in OEIS sequence A194136. Other geometrical investigations into general position sets include [11, 47], as well as a random version in [10, 30, 31]. The general position subset selection problem in discrete geometry also has an extensive literature, including [53, 104], of which we will not attempt to give an exhaustive overview.
@@
Figure 1: A no-three-in-line configuration of 120 pawns on a $60\times 60$ chessboard from [111], reproduced with kind permission from T. Prellberg (with lines connecting the vertices on the right to show the complexity of the problem)
@@ 1.1 Terminology
For a positive integer $k$ we will use the notation $k]=\{1,\ldots,k\}$. We will denote the order and the size of a graph $G$ by $n(G)$ and $m(G)$ respectively. If $u\in V(G)$, then $N_{G}(u)$ denotes the open neighbourhood of $u$ and $N_{G}[u]$ its closed neighbourhood. Vertices $u$ and $v$ of a graph $G$ are true twins if $N_{G}[u]=N_{G}[v]$. A vertex is a support vertex if it is adjacent to a leaf. A vertex $u$ of $G$ is simplicial if its neighbourhood induces a clique. The set of all simplicial vertices of a graph will be denoted by $S(G)$ and their number by $\operatorname{s}(G)$. The number of leaves of a tree $T$ will be written as $\ell(T)$ (so $\ell(T)=\operatorname{s}(T)$). The join of $G$ and $H$ is denoted by $G\vee H$. For $X\subseteq V(G)$, $G[X]$ is the subgraph of $G$ induced by $X$. By $\alpha(G)$ and $\omega(G)$ we represent the independence number and the clique number of $G$ respectively.
@@
If $G$ and $H$ are graphs, then their Cartesian product, direct product, strong product, and lexicographic product are respectively denoted by $G\operatorname{\,\square\,}H$, $G\times H$, $G\boxtimes H$, and $G\circ H$. Each of these products has the vertex set $V(G)\times V(H)$. Vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G\operatorname{\,\square\,}H$ if either $g=g^{\prime}$ and $hh^{\prime}\in E(H)$, or $gg^{\prime}\in E(G)$ and $h=h^{\prime}$. Vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G\times H$ if $gg^{\prime}\in E(G)$ and $hh^{\prime}\in E(H)$. For the strong product we have $E(G\boxtimes H)=E(G\operatorname{\,\square\,}H)\cup E(G\times H)$. Finally, vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G\circ H$ if either $gg^{\prime}\in E(G)$, or $g=g^{\prime}$ and $hh^{\prime}\in E(H)$. If $h\in V(H)$ and $*\in\{\operatorname{\,\square\,},\times$$\boxtimes,\circ\}$, then the subgraph of GHG*H induced by the vertex set $\{(g,h):\ g\in V(G)\}$ is called an $G$-layer of GHG*H and denoted by $G^{h}$. For $g\in V(G)$ the $H$-layer $^{g}H$ is defined analogously.
@@ 2 General Position Number
General position sets have a structure that can be characterised. To do this, we need the following definitions. Let $G$ be a connected graph, $X\subseteq V(G)$, and ${\cal P}=\{X_{1},\ldots,X_{p}\}$ a partition of $X$. Then $\cal P$} is distance-constant if for any $\in[p], iji\neq j$, the distance $d_{G}(u,v)$, $u\in X_{i}$, $v\in X_{j}$, is independent of the selection of $u$ and $v$. If $\cal P$} is a distance-constant partition, and $\in[p], iji\neq j$, then the distance $d_{G}(X_{i},X_{j})$ can be defined as the distance between a vertex of $X_{i}$ and a vertex of $X_{j}$. A distance-constant partition $\cal P$} is in-transitive if $d_{G}(X_{i},X_{k})\neq d_{G}(X_{i},X_{j})+d_{G}(X_{j},X_{k})$ holds for arbitrary pairwise distinct $i,j,k\in[p]$.
@@ 2.1 Bounds
Graphs with $\operatorname{gp}(G)\in\{2,n(G),n(G)-1\}$ are characterised in the paper [25]. In particular, $\operatorname{gp}(G)=n(G)$ if and only if $G=K_{n}$, and $\operatorname{gp}(G)=2$ if and only if $G\in\{C_{4},P_{n}\}$ [25, Theorem 2.10]. By [25, Theorem 2.4], for any connected graph $G$ we have $\operatorname{gp}(G)\leq n(G)-\mathrm{diam}(G)+1$. Consequently, [25, Theorem 3.1] proves that $\operatorname{gp}(G)=n(G)-1$ if and only if $G=K_{1}+\bigcup_{j}m_{j}K_{j}$ with $\sum m_{j}\geq 2$ or $G=K_{n}-\{e_{1},\ldots,e_{k}\}$ with $k\in[n-3]$, where the $e_{i}$’s are edges in $K_{n}$ that are all incident to a common vertex.
@@
[126, Theorem 3.1] If $G$ is a connected graph of order at least four, then ${\rm gp}(G)=n(G)-2$ if and only if $G\in\cup_{i=1}^{8}{\cal F}_{i}.$
@@
[91, Theorem 3.1] (Isometric Cover Lemma) If $\left\{H_{1},\ldots,H_{k}\right\}$ is an isometric cover of $G$, then
@@
The isometric-path number of a graph $G$, which we will write as $\rho(G)$, is the smallest number of isometric paths (shortest paths) needed to cover all the vertices of $G$ (note that $\operatorname{ip}(G)$ is often used for this invariant, but we wish to avoid a notation clash with the independent position number). Similarly, the isometric cycle number of $G$, written as $\rho_{C}(G)$, is the smallest number of isometric cycles required to cover all the vertices of $G$. If $G$ cannot admit a cover by isometric cycles, we set $\rho_{C}(G)=\infty$. Then, the Isometric Cover Lemma yields that for any graph $G$, $\operatorname{gp}(G)\leq 2\rho(G)$, and $\operatorname{gp}(G)\leq 3\rho_{C}(G)$ (see [91, Corollary 3.2]).
@@
[91, Theorem 3.3] If $R$ is a general position set of a graph $G$ and $v\in R$, then
@@
[56, Theorem 4.1] If $G$ is a connected graph, then $\operatorname{gp}(G)\leq\alpha^{\omega}(G)$. Moreover, if $\operatorname{diam}(G)\in[2]$, then $\operatorname{gp}(G)=\alpha^{\omega}(G)$.
@@ 2.1.1 Strong resolving graphs
[81, Theorem 3.1] If $G$ is a connected graph, then $\operatorname{gp}(G)\geq\omega(G\operatorname{{}_{SR}})$. Moreover, equality holds if and only if $G$ contains a $\operatorname{gp}$-set that induces a complete subgraph of $G\operatorname{{}_{SR}}$.
@@
[81, Proposition 2.4] If $G$ has no true twins and $\operatorname{diam}(G)=2$, then $\operatorname{gp}(G)=\omega(G\operatorname{{}_{SR}})$ if and only if $\operatorname{gp}(G)=\alpha(G)$.
On the other hand, as demonstrated in [81, Proposition 3.5], for any integers $r\geq t\geq 2$, there exists a graph $G$ such that $\operatorname{gp}(G)=r$ and $\omega(G\operatorname{{}_{SR}})=t$.
@@ 2.2 Complexity
[91, Theorem 5.1] General position set is NP-Complete.
@@ 2.3 Exact results
As the general position problem is NP-complete, it makes sense to consider the restriction of the problem to some simple graph classes. We recall from the seminal papers [25, 91] the following results. If $n\geq 2$, then $\operatorname{gp}(K_{n})=n$. If $n\geq 5$, then $\operatorname{gp}(C_{n})=3$, while $\operatorname{gp}(C_{4})=2$. If $n\geq 2$, then $\operatorname{gp}(P_{n})=2$. Further, if $G$ is a block graph, then $\operatorname{gp}(G)=\operatorname{s}(G)$ [91, Theorem 3.6]. Hence in particular, if $T$ is a tree, then $\operatorname{gp}(T)=\ell(T)$ [25, Theorem 2.5].
@@ 2.3.1 Kneser graphs
The Kneser graph $K(n,k)$ is the graph whose vertex set consists of all $k$-subsets of the set $[n]$, with an edge between any two subsets if and only if they are disjoint. Ghorbani et al. [56] computed $\operatorname{gp}(K(n,2))$ and $\operatorname{gp}(K(n,3))$ for all $n$. In particular, [56, Theorem 2.2] gives $\operatorname{gp}(K(n,2))=n-1$ when $n\geq 7$, and [56, Theorem 2.4] gives $\operatorname{gp}(K(n,3))=\binom{n-1}{2}$ when $n\geq 9$. In [56, Theorem 2.3], it is proved that for any fixed $k$, if $n$ is sufficiently large, the equality $\operatorname{gp}(K(n,k))=\binom{n-1}{k-1}$ holds. The result reads as follows.
@@
[56, Theorem 2.3] Let $n$ and $k$ be positive integers with $n\geq 3k-1$. If, for all $t$ such that $2\leq t\leq k$, the inequality $k^{t}\binom{n-t}{k-t}+t\leq\binom{n-1}{k-1}$ holds, then $\operatorname{gp}(K(n,k))=\binom{n-1}{k-1}$.
For fixed $k$ and $t=2$, the above inequality is satisfied when $n\geq k^{3}-k^{2}+2k-1$, while Patkós [102, Theorem 1.2] improved this result by proving that the same conclusion holds for $n\geq 2.5k-0.5$. The result is stated as follows.
@@
[102, Theorem 1.2] If $n,k\geq 4$ are integers with $n\geq 2k+1$, then $\operatorname{gp}(K(n,k))\leq\binom{n-1}{k-1}$ holds. Moreover, if $n\geq 2.5k-0.5$, then we have $\operatorname{gp}(K(n,k))=\binom{n-1}{k-1}$, while if $2k+1\leq n<2.5k-0.5$, then $\operatorname{gp}(K(n,k))<\binom{n-1}{k-1}$ holds.